A given amount of oxygen gas has a volume of 25.0 L at a temperature of 37oC and a pressure of 1.0 atm. At what temperature would this gas occupy a volume of 22.0 L at a pressure of 1.0 atm?


1
Expert's answer
2021-04-07T11:38:01-0400

P1V1/T1 = P2V2/T2

P1 = 1 atm

V1 = 25 L

T1 = 37+273 = 310K

P2 = 1 atm

V2 = 22 L

T2 = ?

= 1×25/310 = 21×1/X

= 0.080645 = 21/X

0.080645X = 21

X = 21/0.080645 = 260.4K

= 260.4 - 273= -12.6

= -12.6°C